Twice the sum of a number and eight, equals three times the sum of the number and four. Find the number?
I don't understand what it is asking..

Answer by ankor@dixie-net.com(22740)

Let n = "the number"
:
Write an equation for what it says:
:
Twice the sum of a number and eight, equals three times the sum of the number and four.
2(n + 8) = 3(n + 4)
Multiply what's inside brackets
2n + 16 = 3n + 12
:
16 - 12 = 3n - 2n
:
4 = n
;
the number is 4:
:
:
Check this by replacing n with 4 in the original equation:
2(4 + 8) = 3(4 + 4)
2(12) = 3(8)
:
Did I make this understandable to you?